# Asymptotic stability of composite waves of two viscous shocks for relaxed compressible Navier-Stokes equations

### v2

@@ 4.1. $L^{2}$ energy estimates
$|B_{6}|+|B_{7}|$ $\leq C\varepsilon_{1}\sum\limits_{i=1}\limits^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-(\widetilde{v}_{i})^{X_{i}}_{x}|\right\|_{L^{2}}+\frac{1}{64}G+C\sum\limits_{i=1}\limits^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-(\widetilde{v}_{i})^{X_{i}}_{x}|\right\|_{L^{2}}^{2}$
$\leq\frac{1}{64}G+C\varepsilon_{1}\delta_{1}\delta_{2}(\delta_{1}^{\frac{1}{2}}+\delta_{2}^{\frac{1}{2}})e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\delta_{1}^{2}\delta_{2}^{2}(\delta_{1}+\delta_{2})e^{-C\min\{\delta_{1},\delta_{2}\}t}$
$\leq\frac{1}{64}G+C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Combining all the above estimates, we derive that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\frac{1}{2}G_{1}+G_{3}+\frac{1}{4}G+\frac{1}{2}\mathcal{G}^{s}$
$\leq\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+C_{2}\mathcal{D}.$
@@
$\int_{\mathbb{R}}a(t,x)\eta(U(t,x)|\widetilde{U}(t,x))\mathrm{d}x+\int_{0}^{t}\left(\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}+G_{1}+G_{3}+G+\mathcal{G}^{s}\right)\mathrm{d}t$
$\leq C\int_{\mathbb{R}}a(0,x)\eta(U(0,x)|\widetilde{U}(0,x))\mathrm{d}x+C\delta_{0}+C_{2}\int_{0}^{t}\mathcal{D}\mathrm{d}t.$
@@ 4.2. High-order energy estimates
$\qquad+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}(U)\mathrm{d}t+C(\delta_{1}+\delta_{2}+\varepsilon_{1})\int_{0}^{t}\left(\|\left(\partial_{x}\Phi,\partial_{x}\Psi\right)\|_{H^{1}}^{2}+G(U)\right)\mathrm{d}t$
$\qquad+C\int_{0}^{t}\sum\limits_{i=1}^{2}\delta_{i}^{2}|\dot{X}_{i}|^{2}\mathrm{d}t+C\delta_{0},$
@@
Multiplying the above equations by $-p^{\prime}(v)\partial_{x}^{k}\Phi,\partial_{x}^{k}\Psi,\frac{1}{\mu}\partial_{x}^{k}Q$, respectively, and integrating over $\mathbb{R}$, we get
@@
$\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||v-\widetilde{v}||(\widetilde{v}_{i})^{X_{i}}_{x}|\mathrm{d}x\leq C\delta_{i}\left(\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||v-\widetilde{v}|^{2}\mathrm{d}x\right).$
@@
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}.$
@@
$R_{6}^{1}$ $\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x$
$+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}+C\varepsilon_{1}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
@@
$\leq C(\varepsilon_{1}^{2}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|\partial_{xx}\Psi|^{2}\mathrm{d}x\right)+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{xx}\Phi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}$
@@
$\qquad+C(\delta_{1}+\delta_{2})G^{s}+C\varepsilon_{1}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
@@
$v(\Pi-\widetilde{\Pi})_{x}-(v\Pi-\widetilde{v}\widetilde{\Pi})_{x}=-(v-\widetilde{v})\widetilde{\Pi}_{x}-(v-\widetilde{v})_{x}(\Pi-\widetilde{\Pi})-(v-\widetilde{v})_{x}\widetilde{\Pi}-\widetilde{v}_{x}(\Pi-\widetilde{\Pi}),$
@@
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x+G\right)+(\delta_{1}+\delta_{2})\int_{\mathbb{R}}(\partial_{x}\Phi)^{2}\mathrm{d}x+(\delta_{1}+\delta_{2})G^{s}.$
@@
$R_{7}^{1}\leq\frac{1}{8}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x+(\varepsilon_{1}+\delta_{1}+\delta_{2})G+(\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}(\partial_{x}\Phi)^{2}\mathrm{d}x+G^{s}\right)+C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
@@
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left\|\sqrt{\frac{v}{\mu}}Q\right\|_{H^{2}}+(\delta_{1}+\delta_{2})\left\|\partial_{x}\Phi\right\|_{H^{1}}+(\delta_{1}+\delta_{2})G^{s}.$
@@
$R_{7}^{2}$ $\leq\frac{1}{8}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{xx}Q)^{2}\mathrm{d}x$
$+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left\|\sqrt{\frac{v}{\mu}}Q\right\|_{H^{1}}+C(\delta_{1}+\delta_{2})\left(G^{s}+\left\|\partial_{x}\Phi\right\|_{H^{1}}\right)+C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
@@ 4.3. Dissipative estimates
Under the hypotheses of Proposition 3.1, there exist $C,C_{1},\nu>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$, we have
$\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu\|\Psi\|_{H^{1}}^{2}+C(\nu)\|\partial_{x}\Phi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+C\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t$ (4.16)
$+C\left(\|\Psi_{0}\|_{H^{1}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}\right)+C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}(U)\mathrm{d}t+C\delta_{0},$
where $G^{s}(U)$ is defined in (3.7).
@@
Multiplying the equation (4.14)2italic-(4.14subscriptitalic-)$_{2}$ by $\partial_{x}^{k+1}\Phi$ for $k=0,1,$ and integrating over $(0,t)\times\mathbb{R}$, we get
@@
$-\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}\Psi\partial_{x}^{k+2}\Psi\mathrm{d}x\mathrm{d}t=\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t,$
@@
$M_{1}^{k}\leq$ $\nu\int_{\mathbb{R}}\left(\partial_{x}^{k}\Psi(t)\right)^{2}\mathrm{d}x+C(\nu)\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi(t)\right)^{2}\mathrm{d}x+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t$
@@
$M_{2}^{k}\leq C\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+\delta_{i}\int_{0}^{t}\int_{\mathbb{R}}(\partial_{x}^{k+1}\Phi)^{2}\mathrm{d}x\mathrm{d}t\right)$
@@
$M_{3}^{k}\leq\frac{1}{2}\int_{0}^{t}\int_{\mathbb{R}}-p^{\prime}(v)\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t.$
@@
$\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu_{1}\tau\|Q\|_{H^{1}}^{2}+C(\nu_{1})\|\partial_{x}\Psi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+\frac{1}{2C_{1}}\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t$ (4.17)
$+C\left(\tau\|Q_{0}\|_{H^{1}}^{2}+\|\partial_{x}\Psi_{0}\|_{H^{1}}^{2}\right)+C\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{0}^{t}(G^{s}(U)+G(U))\mathrm{d}t+C\delta_{0},$
where $C_{1}$ is the constant in Lemma 4.7, and $G^{s}(U),G(U)$ are defined in (3.7) and (4.4), respectively.
@@
Multiplying the equation (4.14)3italic-(4.14subscriptitalic-)$_{3}$ by $\partial_{x}^{k+1}\Psi$ for $k=0,1$, and integrating over $(0,t)\times\mathbb{R}$, we get
@@
Firstly, doing integration by part and using equation (4.14)2italic-(4.14subscriptitalic-)$_{2}$, we get

### v3

@@ 4.1. $L^{2}$ energy estimates
$|B_{6}|+|B_{7}|$ $\leq C\varepsilon_{1}\sum\limits_{i=1}\limits^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-(\widetilde{v}_{i})^{X_{i}}_{x}|\right\|_{L^{2}}+\kappa G+C\sum\limits_{i=1}\limits^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-(\widetilde{v}_{i})^{X_{i}}_{x}|\right\|_{L^{2}}^{2}$
$\leq\kappa G+C\varepsilon_{1}\delta_{1}\delta_{2}(\delta_{1}^{\frac{1}{2}}+\delta_{2}^{\frac{1}{2}})e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\delta_{1}^{2}\delta_{2}^{2}(\delta_{1}+\delta_{2})e^{-C\min\{\delta_{1},\delta_{2}\}t}$
$\leq\kappa G+C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Noting that $\frac{1}{t^{2}}$ is not integrable near $t=0$, so we first to get the estimate for $t\geq 1$. Then, for any $\delta_{1},\delta_{2}\in(0,\delta_{0})$, combining all the above estimates, we derive that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\frac{1}{2}G_{1}+G_{3}+(1-\kappa-C\delta^{\frac{1}{2}}_{0})G+C_{1}(1-C(\delta_{0}+\varepsilon_{1}))\mathcal{G}^{s}$
$\leq C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\delta_{i}e^{-C\delta_{i}t}+\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+\frac{1}{t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x$
$\quad+\frac{3\mu}{4v_{m}}(1+\kappa+C\varepsilon_{1})\int_{\mathbb{R}}|p^{\prime}(v)||\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x.$
By using Grönwall inequality, it holds
$\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\int_{1}^{t}\left(\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\frac{1}{2}G_{1}+G_{3}+(1-\kappa-C\delta^{\frac{1}{2}}_{0})G+C_{1}(1-C(\delta_{0}+\varepsilon_{1}))\mathcal{G}^{s}\right)\mathrm{d}t$
$\leq C\left(\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x\Big|_{t=1}+\frac{\delta_{1}\delta_{2}}{\min\{\delta_{1},\delta_{2}\}}\right)\times\exp\left(\int_{1}^{t}\Big(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\delta_{i}e^{-C\delta_{i}s}+\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}s}+\frac{1}{s^{2}}\Big)\mathrm{d}s\right)$
$\quad+\frac{3\mu}{4v_{m}}(1+\kappa)\int_{1}^{t}\int_{\mathbb{R}}|p^{\prime}(v)||\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x\mathrm{d}s$
$\leq C\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x\Big|_{t=1}+\frac{3\mu}{4v_{m}}(1+\kappa+C\varepsilon_{1})\int_{1}^{t}\int_{\mathbb{R}}|p^{\prime}(v)||\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x\mathrm{d}s+C\delta_{0}.$
On the other hand, firstly, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x=-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{M}|\dot{X}_{i}|^{2}+\sum\limits_{i=1}\limits^{2}\left(\dot{X}_{i}\sum\limits_{j=3}\limits^{8}Y_{ij}\right)+\sum\limits_{i=1}\limits^{7}B_{i}-\sum\limits_{i=1}\limits^{3}G_{i}-G.$
Using Young’s inequality, we get
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\sum\limits_{i=1}\limits^{3}G_{i}+G+C_{1}\mathcal{G}^{s}\leq\sum\limits_{i=1}\limits^{2}\left(\frac{C}{\delta_{i}}\sum\limits_{j=3}\limits^{8}|Y_{ij}|^{2}\right)+\sum\limits_{i=1}\limits^{7}B_{i}+C_{1}\mathcal{G}^{s}.$
By the Hölder inequality, Lemma 2.1, we have
$\sum\limits_{j=3}\limits^{8}|Y_{ij}|$ $\leq C\|(\widetilde{v}_{i})^{X_{i}}_{x}\|_{L^{2}}\left(\|p(v)-p(\widetilde{v})\|_{L^{2}}+\|u-\widetilde{u}\|_{L^{2}}+\sqrt{\tau}\|S-\widetilde{S}\|_{L^{2}}\right)$
$+C\|(a_{i})^{X_{i}}_{x}\|_{L^{\infty}}\left(\|p(v)-p(\widetilde{v})\|_{L^{2}}^{2}+\|u-\widetilde{u}\|_{L^{2}}^{2}+\tau\|S-\widetilde{S}\|_{L^{2}}^{2}\right)$
$\leq C\delta_{0}\varepsilon_{1},$
which yields
$\sum\limits_{i=1}\limits^{2}\left(\frac{C}{\delta_{i}}\sum\limits_{j=3}\limits^{8}|Y_{ij}|^{2}\right)\leq C\varepsilon_{1}^{2}\delta_{0}.$
Similarly, we have
$\sum\limits_{i=1}\limits^{7}B_{i}\leq C\left(\|(a_{i})^{X_{i}}_{x}\|_{L^{\infty}}+\|\widetilde{S}\|_{L^{\infty}}\right)\|p(v)-p(\widetilde{v})\|_{L^{2}}^{2}+C(\delta_{1}+\delta_{2})G\leq C\varepsilon_{1}^{2}\delta_{0}+C\delta_{0}G,$
and
$\mathcal{G}^{s}\leq C\sum\limits_{i=1}\limits^{2}\|(\widetilde{v}_{i})^{X_{i}}_{x}\|_{L^{\infty}}\|p(v)-p(\widetilde{v})\|_{L^{2}}^{2}\leq C\varepsilon_{1}^{2}\delta_{0}.$
Hence, combining above estimates, we derive that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+G_{1}+G_{3}+(1-C\delta_{0})G+C_{1}\mathcal{G}^{s}\leq C\delta_{0}.$
Then, for a short time $0\leq t\leq 1$, we have
$\sup_{0\leq t\leq 1}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\int_{0}^{1}\left(\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+G_{1}+G_{3}+(1-C\delta_{0})G+C_{1}\mathcal{G}^{s}\right)\mathrm{d}t\leq\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x\Big|_{t=0}+C\delta_{0}.$
@@
$\int_{\mathbb{R}}a(t,x)\eta(U(t,x)|\widetilde{U}(t,x))\mathrm{d}x+\int_{0}^{t}\left(\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\frac{1}{2}G_{1}+G_{3}+(1-\kappa-C\delta^{\frac{1}{2}}_{0})G+C_{1}(1-C(\delta_{0}+\varepsilon_{1}))\mathcal{G}^{s}\right)\mathrm{d}t$
$\leq C\int_{\mathbb{R}}a(0,x)\eta(U(0,x)|\widetilde{U}(0,x))\mathrm{d}x+\frac{3\mu}{4v_{m}}(1+\kappa)\int_{0}^{t}\int_{\mathbb{R}}|p^{\prime}(v)||\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x\mathrm{d}s+C\delta_{0}.$
@@ 4.2. High-order energy estimates
$+C\delta_{0}\int_{0}^{t}G^{s}(U)\mathrm{d}t+C(\delta_{0}+\varepsilon_{1})\int_{0}^{t}\left(\|\left(\partial_{x}\Phi,\partial_{x}\Psi\right)\|_{H^{1}}^{2}+G(U)\right)\mathrm{d}t+C\int_{0}^{t}\sum\limits_{i=1}^{2}\delta_{i}^{2}|\dot{X}_{i}|^{2}\mathrm{d}t+C\delta_{0},$
@@
Multiplying the above equations by $p^{\prime}(v)\partial_{x}^{k}\Phi,\partial_{x}^{k}\Psi,\frac{1}{\mu}\partial_{x}^{k}Q$, respectively, and integrating over $\mathbb{R}$, we get
@@
$\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||v-\widetilde{v}||(\widetilde{v}_{i})^{X_{i}}_{x}|\mathrm{d}x$
$\leq C\delta_{i}\left(\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||v-\widetilde{v}|^{2}\mathrm{d}x\right)$
$\leq C\delta_{i}\left(\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||\phi_{i}(v-\widetilde{v})|^{2}\mathrm{d}x+\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|(1-\phi_{i}^{2})|v-\widetilde{v}|^{2}\mathrm{d}x\right)$
$\leq C\delta_{i}\left(\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+G^{s}(U)+\varepsilon_{1}\delta_{i}^{2}e^{-C\delta_{i}t}\right).$
@@
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}.$
@@
$R_{6}^{1}\leq$ $C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}$
$+C\varepsilon_{1}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}.$
@@
$\leq C(\varepsilon_{1}^{2}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|\partial_{xx}\Psi|^{2}\mathrm{d}x\right)+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{xx}\Phi|^{2}\mathrm{d}x$
$\quad+C(\delta_{1}+\delta_{2})G^{s}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}$
@@
$\qquad+C(\delta_{1}+\delta_{2})G^{s}+C\varepsilon_{1}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}.$
@@
[…]
@@
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x+G\right)+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}(\partial_{x}\Phi)^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}.$
@@
$R_{7}^{1}\leq\frac{1}{8}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x+C(\varepsilon_{1}+\delta_{1}+\delta_{2})G+C(\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}(\partial_{x}\Phi)^{2}\mathrm{d}x+G^{s}\right)$
$+C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}.$
@@
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left\|\sqrt{\frac{v}{\mu}}Q\right\|_{H^{2}}+(\delta_{1}+\delta_{2})\left\|\partial_{x}\Phi\right\|_{H^{1}}+(\delta_{1}+\delta_{2})G^{s}+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}.$
@@
$R_{7}^{2}$ $\leq\frac{1}{8}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{xx}Q)^{2}\mathrm{d}x+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left\|\sqrt{\frac{v}{\mu}}Q\right\|_{H^{1}}+C(\delta_{1}+\delta_{2})\left(G^{s}+\left\|\partial_{x}\Phi\right\|_{H^{1}}\right)$
$+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\delta_{i}^{3}e^{-C\delta_{i}t}+C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
@@ 4.3. Dissipative estimates
Under the hypotheses of Proposition 3.1, there exist $C,\nu>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$, we have
$(1-\kappa-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{t}\|\sqrt{|p^{\prime}(v)|}\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t$ (4.18)
$\leq\nu\|\Psi\|_{H^{1}}^{2}+C(\nu)\|\partial_{x}\Phi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+C\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t$
$+C\left(\|\Psi_{0}\|_{H^{1}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}\right)+(1+C\delta_{0})\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t+C\delta_{0}\int_{0}^{t}G^{s}(U)\mathrm{d}t+C\delta_{0},$
where $\kappa$ is small constant to be determined later and $G^{s}(U)$ is defined in (3.7).
@@
[…]
@@
$\int_{\mathbb{R}}\partial_{x}^{k}\Psi(t)\partial_{x}^{k+1}\Phi(t)\mathrm{d}x\leq\nu\int_{\mathbb{R}}(\partial_{x}^{k}\Psi(t))^{2}\mathrm{d}x+C(\nu)\int_{\mathbb{R}}(\partial_{x}^{k+1}\Phi(t))^{2}\mathrm{d}x,$
$\int_{\mathbb{R}}\partial_{x}^{k}\Psi_{0}\partial_{x}^{k+1}\Phi_{0}\mathrm{d}x\leq C\int_{\mathbb{R}}(\partial_{x}^{k}\Psi_{0})^{2}\mathrm{d}x+C\int_{\mathbb{R}}(\partial_{x}^{k+1}\Phi_{0})^{2}\mathrm{d}x,$
@@
$\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}\Psi\partial_{x}^{k+2}\Psi\mathrm{d}x\mathrm{d}t=\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t,$
@@
$M_{1}^{k}\leq$ $\nu\int_{\mathbb{R}}\left(\partial_{x}^{k}\Psi(t)\right)^{2}\mathrm{d}x+C(\nu)\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi(t)\right)^{2}\mathrm{d}x+(1+C\delta_{0})\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t$
@@
$M_{2}^{k}\leq C\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+\delta_{i}\int_{0}^{t}\int_{\mathbb{R}}|p^{\prime}(v)|(\partial_{x}^{k+1}\Phi)^{2}\mathrm{d}x\mathrm{d}t\right)$
@@
$M_{3}^{k}\leq\kappa\int_{0}^{t}\int_{\mathbb{R}}-p^{\prime}(v)\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t.$
@@
$\frac{\mu}{2}(1-C\delta_{0})\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu_{1}\tau\|Q\|_{H^{1}}^{2}+C(\nu_{1})\|\partial_{x}\Psi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t$ (4.19)
$\quad+\mu(\kappa+C\delta_{0})\int_{0}^{t}\|\sqrt{|p^{\prime}(v)|}\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+C\left(\tau\|Q_{0}\|_{H^{1}}^{2}+\|\partial_{x}\Psi_{0}\|_{H^{1}}^{2}\right)$
$\quad+C\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t+C(\varepsilon_{1}+\delta_{0})\int_{0}^{t}G^{s}(U)\mathrm{d}t+\frac{v_{m}+C(\delta_{0}+\varepsilon_{1})}{2}\int_{0}^{t}G(U)\mathrm{d}t+C\delta_{0},$
where $\kappa$ is small constant to be determined later, $G^{s}(U),G(U)$ are defined in (3.7) and (4.4), respectively.
@@
[…]
@@
[…]